Hi please help me solve this equation:Solving Linear Systems-Solve the linear system using the substitution method.
x=17-4y
y=x-2

Answer by boilpoil(127)

x=17-4y
y=x-2

x=17-4(x-2) ---Substitute the value of y into the first equation
x=17-4x+8
5x=25
x=5

y=x-2 ---Substitute the value of x into the second equation
y=5-2
y=3